### v1

@@ III-B2 Architecture
A detailed description of the model architecture is provided in Table II.
TABLE II: U-Net architecture.
@@ III-B3 Training
For training the U-Net, the objective was to recreate Otsu-thresholded patches of $320\times 320$ resolution based on the original image patches augmented with synthetic noise. At each epoch, 16 patches were randomly cropped from the well-defined sampleable region of each image. The patches were further augmented by applying random rotations of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$ or $270^{\circ}$. The network was implemented in Python using the PyTorch framework [^1] https://pytorch.org/. It was trained with a batch size of 32 and a learning rate of $10^{-3}$. The loss function used was Binary Cross Entropy with Logits Loss (BCEWithLogitsLoss), which internally applies a sigmoid function, allowing for training with the segmentation targets in the [0,1] range. The model was trained for a total of 5 epochs with Adam optimizer. Training the model for longer caused overfitting, reducing the segmentation accuracy under real degradations.
Training was conducted using a single NVidia A100 GPU in the Google Colab environment222colab.research.google.com. Error cases in segmentation are discussed in Appendix VI-B.
@@ III-C Border and inner paths
The topological skeleton, also known as the medial axis, was generated using the skimage implementation of medial axis skeletonization. It represents the set of points equidistant to at least two points on the boundary using the Euclidean distance transform. Within the magnetic stripes, it forms the central axis along the dark patterns, but includes spurious branches caused by minor variations in the border shape. To accurately model the inner paths within the dark patterns, these spurious branches were removed. This was done by matching the terminal points of the skeleton with the terminals detected by TM-CNN defect detection [19]. Terminals that did not match were trimmed back to a junction point, provided their length did not exceed 45 pixels.
@@
The boundary of the stripe patterns were produced through contour finding using OpenCV’s findContours, which is based on the Suzuki-Abe algorithm [60]. Fig. 3(c, f) showcases the boundary and skeleton images respectively.
@@ III-D Graph mapping
To process the skeleton, terminal coordinates are initially pushed onto a stack. We then use each coordinate as a starting point (kernel) to trace the trimmed skeleton using 8-connectivity on a temporary image copy. Pixels are marked as 255 upon visitation to avoid redundant processing. If a trace reaches another defect, we verify its representation as a graph node and create an edge connecting it to the starting kernel node, saving the path taken. When a junction is encountered —a pixel with several unvisited neighbors— a junction is created. To represent the junction, a node is created for each of the neighboring pixels, together constituting a single junction in the graph. All unvisited neighbors are then pushed onto the stack so that every branch can be traced.
@@
3:while NotEmpty(defect_coordsdefect\_coordsdo
@@ III-E Length measurements
Smooth splines were generated through an interpolation process, sampling from the pixel paths using the splprep function from the SciPy library  [^3] https://scipy.org/. The degree of smoothing is determined by the coefficient $s$ which limits the total square error as shown in equation:
@@
where $g(x_{j})$ is the smoothed spline and $y_{j}$ are pixel samples from the path. The value of $s$ was chosen as equal to the length of the path in pixels, to ensure that paths of different lengths were smoothed to a similar degree [61]. Points on the spline were then sampled using splev, also from SciPy, with the number of samples being twice the number of pixels in the original path. To generate the Euclidean length of each path, the spline was numerically integrated based on the sampled points.
@@ IV-A Dark Area
From the U-Net segmented images, the area occupied by the dark region was estimated (Fig. 9). Measurements from the positive annealing protocol are marked with “+”, while those from the negative protocol are marked with “-”. Results for both protocols were averaged over six trials each, with the standard deviation shown as a shaded region. An interleaved pattern emerged in both the positive and negative sequences, depending on whether the measurement is Type A (plotted in orange), where the applied field is aligned with the magnetization of the black region, or Type B (plotted in blue), where it is aligned with the magnetization of the white region.
@@
Magnetic moments couple to the external magnetic field via Zeeman coupling and energetically favor alignment with the field. During imaging, the magnetic field is zero, and thus an equal distribution of dark and bright areas is expected. Meanwhile, the annealing protocol involves a magnetic field quench and the majority of magnetic moments align with the field direction before the quench. This is particularly pronounced in the early stages, where fields above the saturation field are applied, driving the YIG film into a fully-polarized state. Due to the memory effect of the system, an imbalance between the dark and bright domain areas is observed. For instance, in Type A measurements, a larger dark domain area is observed compared to Type B.
@@ IV-B Defect Count
Fig. 10 shows the evolution of defect counts during the annealing procedure. In Type B measurements, the quenched state begins with a higher number of defects, which gradually decreases and stabilizes at a lower value as annealing progresses, consistent with Ref.[19]. In contrast, Type A measurements start with a slightly lower defect count in the quenched state, which stabilizes at a higher value in the annealed state, eventually converging toward the defect count of Type B. Although the step dependence differs between Type A and Type B, junctions and terminals within the same sequence display similar step-dependent trends. This behavior reflects the topological nature of these defects: junctions and terminals cannot be created or annihilated individually through smooth stripe deformations, but only through pair creation or annihilation [20].
Annealing is generally expected to increase coherence and reduce the number of defects. At first glance, the results in Fig. 10 seem contradictory. Yet, in our measurements, the defect count does not monotonically decrease during the protocol. This apparent inconsistency is resolved by considering the change in stripe period reported in Ref. [20]. During annealing, the stripe period decreases by roughly 20% between steps 12 and 24. Since defects are tied to stripe structures, a shorter period naturally supports a higher defect density. Indeed, for Type A measurements, the defect count normalized by the stripe period decreases [20]. Since the defect count in Type B measurements in the annealed state is larger than that in Type A, the normalized defect count in Type B likewise decreases.
Comparison of Type A and Type B measurements shows that before step 12 the defect count of Type B is substantially higher, while beyond this step the difference between the two becomes negligible. The qualitative differences in defect counts arise from distinct dominant dynamics at different stages: in the early steps, local defect dynamics prevail, while in the later steps the variation in stripe period becomes the primary factor, as previously reported [20].
@@
Figure 10: Number of defects averaged over six trials. Green illustrate the count of junctions, and blue show the count of terminals. Type A measurements are depicted in lighter colors, and Type B in darker colors. Shaded regions represent the standard deviation.
@@ IV-C Medial axis length between defects
The second factor is the correlation between defects. In Ref. [20], it was reported that strong correlations develop between junctions and terminals, and around step 12, where the system undergoes a transition from the quenched to the annealed state, the annihilation of junction-terminal pairs contributes to an increase in the coherence of the system. Since short junction–terminal pairs are preferentially annihilated, the mean length of the remaining pairs can increase. This trend is especially pronounced in Type B measurements for junction–terminal pairs.
Figure 11: Mean medial axis lengths between topological defects averaged over six trials. Green segments show the length between adjacent junctions, and yellow the length between adjacent junctions and terminals. Lighter shades denote Type A measurements, and darker shades denote Type B.
@@ IV-D Curvature
Fig. 12(a) displays the border curvature of the labyrinthine pattern. Both Type A and Type B images show similar curvature patterns, though Type A has slightly higher curvature in the quenched state. For both measurements, the curvature peaks around steps 9-11, coinciding with the transition from the quenched to the annealed state.
@@ IV-E Total Length
Fig. 13 (a)-(b) illustrates the total inner and border lengths of the magnetic stripe pattern images. The graphs show a lower average length at the beginning of the annealing protocol (in the quenched state), followed by stabilization at a higher average length in the later steps (in the annealed state). This increase in total length throughout the annealing protocol can be attributed to a more efficient packaging of stripes in the annealed state due to its higher spatial order as well as the reduction of the stripe period [20].

### v2

@@ III-B2 Architecture
A detailed description of the model architecture is provided in Table III.
TABLE III: U-Net architecture.
@@ III-B3 Training
The U-Net was trained to reconstruct Otsu-thresholded binary masks from clear patches with $320{\times}320$ resolution. These patches were corrupted with synthetic noise to simulate degraded regions. Trials $t\in\{1,2,3,4\}$, comprising 296 images, were used for training and validation. The remaining 148 images from trials $t\in\{5,6\}$ were reserved for the test set.
The training set comprised patches from high-local-variance regions, while validation used areas with a local variance under 70% of the overall image variance, ensuring that most patches exhibited substantial degradation. We sampled and manually annotated 64 patches ($320{\times}320$). Performance metrics were calculated without augmentation.
For each training epoch, four patches were randomly sampled from the well-defined region of each training image. The patches were further augmented by applying random rotations of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$ or $270^{\circ}$. The network was implemented in Python using the PyTorch framework [^1] https://pytorch.org/. It was trained with a batch size of 32 and a learning rate of $10^{-3}$. The loss function was Binary Cross Entropy with Logits Loss (BCEWithLogitsLoss), which internally applies a sigmoid function, allowing for training with the segmentation targets in the $[0,1]$ range. The model was trained for a total of 30 epochs with the Adam optimizer. The epoch with the best validation IoU was used for test evaluation.
The resulting loss curves and validation performance metrics are compiled in Appendix B, alongside a quantitative ablation study on the impact of synthetic degradations during training on the model’s performance.
Training was conducted using a single NVIDIA A100 GPU in the Google Colab environment222colab.research.google.com.
III-B4 Testing
The U-Net’s performance was tested for segmenting degraded regions using the held-out testing trials $t\in\{5,6\}$. Analogous to the validation procedure, 64 patches of $320{\times}320$ pixels were randomly sampled from the low-variance areas of the images and manually segmented to serve as the ground truth for evaluation.
Given the recent popularity in Transformer-based segmentation models, we also finetuned a SegFormer-B3 model initialized from ImageNet-1k weights. The model was trained on the same dataset and noise augmentations as the U-Net model.
[…]
TABLE IV: Segmentation performance metrics.
Model F1 score ($\uparrow$) IoU ($\uparrow$) ASSD ($\downarrow$) $\DeltaInner Δ\DeltaBorder$
Otsu 0.948 0.901 1.08 1.38% 2.90%
SegFormer-B3 0.953 0.911 1.01 -0.41% -0.05%
U-Net 0.953 0.910 1.01 -0.15% 1.16%
@@ III-C Border and inner paths
The topological skeleton, also known as the medial axis, was generated using the skimage implementation of medial axis skeletonization. It represents the set of points equidistant to at least two points on the boundary using the Euclidean distance transform. Within the magnetic stripes, it forms the central axis along the dark patterns, but includes spurious branches caused by minor variations in the border shape. To accurately model the inner paths within the dark patterns, these spurious branches were removed. This was done by matching the terminal points of the skeleton with the terminals detected by TM-CNN defect detection [okubo2024]. Terminals that did not match were trimmed back to a junction point, provided their length did not exceed 45 pixels.
@@
The boundary of the stripe patterns were produced through contour finding using OpenCV’s findContours, which is based on the Suzuki-Abe algorithm [suzuki1985]. Fig. 3(c, f) showcases the boundary and skeleton images respectively.
@@ III-D Graph mapping
To process the skeleton, terminal coordinates are initially pushed onto a stack. We then use each coordinate as a starting point (kernel) to trace the trimmed skeleton using 8-connectivity on a temporary image copy. Pixels are marked as 255 upon visitation to avoid redundant processing. If a trace reaches another defect, we verify its representation as a graph node and create an edge connecting it to the starting kernel node, saving the path taken. When a junction is encountered —a pixel with several unvisited neighbors— a junction is created. Because junctions in the pixelated skeleton consist of pixel clusters rather than single points, a node is initialized for each constituent pixel within the junction. These nodes are connected to one another through edges with empty paths to represent a single topological structure. Only the edges connecting the junction to outgoing branches store the corresponding segment path. In this way, the nodes represent a single junction, preserving branch connectivity while avoiding ambiguity during path tracing. The unvisited neighbors are then pushed onto the stack so that every branch can be traced.
@@
3:while NotEmpty(defect_coordsdefect\_coords) do
@@ III-E Length measurements
Smooth splines were generated through an interpolation process, sampling from the pixel paths using the splprep function from the SciPy library. The degree of smoothing is determined by the coefficient $s$ which limits the total square error as shown in equation:
@@
where $g(x_{j})$ is the smoothed spline and $y_{j}$ are pixel samples from the path. The value of $s$ was chosen as equal to the length of the path in pixels, to ensure that paths of different lengths were smoothed to a similar degree [scipy_smoothing_splines]. Points on the spline were then sampled using splev, also from SciPy, with the number of samples being twice the number of pixels in the original path. To generate the Euclidean length of each path, the spline was numerically integrated based on the sampled points.
@@ IV-A Dark Area
From the U-Net segmented images, the area occupied by the dark region was estimated (Fig. 9). Measurements from the positive annealing protocol ($\sigma\in\{+\}$) are marked with “+”, while those from the negative protocol ($\sigma\in\{-\}$) are marked with “-”. Results for both protocols were averaged over six trials each, with the standard deviation shown as a shaded region. An interleaved pattern emerged in both the positive and negative sequences, depending on whether the measurement is Type A (plotted in orange), where the applied field is aligned with the magnetization of the black region, or Type B (plotted in blue), where it is aligned with the magnetization of the white region.
@@
The area of a dark region—corresponding to a ferromagnetic domain—is governed by the interplay between the external magnetic field and competing spin interactions. In both quench and annealing protocols, the applied field plays a central role in shaping domain morphology through the Zeeman coupling, which energetically favors alignment of magnetic moments with the field. Another key contribution in thin-film systems is the magnetic anisotropy energy, which tends to align magnetic moments perpendicular to the film plane. The emergence of stripe patterns in such thin films reflects a competition between interactions at different length scales. Short-range exchange favors uniform spin alignment and thus large domains, while long-range dipole–dipole interactions penalize net out-of-plane magnetization, promoting alternating up- and down-magnetized regions. This frustration stabilizes a modulated stripe phase, with a characteristic width set by the balance between exchange, which favors broader domains, and dipolar interactions, which favor finer alternation.
During imaging, the magnetic field is zero, and thus an equal distribution of dark and bright areas is expected. Meanwhile, the annealing protocol involves a magnetic field quench and the majority of magnetic moments align with the field direction before the quench. This is particularly pronounced in the early stages, where fields above the saturation field are applied, driving the YIG film into a fully-polarized state. Due to the memory effect of the system, an imbalance between the dark and bright domain areas is observed. For instance, in Type A measurements, a larger dark domain area is observed compared to Type B.
@@ IV-B Defect Count
Fig. 10 shows the evolution of defect counts during the annealing procedure. In Type B measurements, the quenched state begins with a higher number of defects, which gradually decreases and stabilizes at a lower value as annealing progresses, consistent with Ref.[okubo2024]. In contrast, Type A measurements start with a slightly lower defect count in the quenched state, which stabilizes at a higher value in the annealed state, eventually converging toward the defect count of Type B. Although the step dependence differs between Type A and Type B, junctions and terminals within the same sequence display similar step-dependent trends. This behavior reflects the topological nature of these defects: junctions and terminals cannot be created or annihilated individually through smooth stripe deformations, but only through pair creation or annihilation [shimizu2023].
Annealing is generally expected to increase coherence and reduce the number of defects. At first glance, the results in Fig. 10 seem contradictory. Yet, in our measurements, the defect count does not monotonically decrease during the protocol. This apparent inconsistency is resolved by considering the change in stripe period reported in Ref. [shimizu2023]. During annealing, the stripe period decreases by roughly 20% between steps 12 and 24. Since defects are tied to stripe structures, a shorter period naturally supports a higher defect density. Indeed, for Type A measurements, the defect count normalized by the stripe period decreases [shimizu2023]. Since the defect count in Type B measurements in the annealed state is larger than that in Type A, the normalized defect count in Type B likewise decreases.
Comparison of Type A and Type B measurements shows that before step 12 the defect count of Type B is substantially higher, while beyond this step the difference between the two becomes negligible. The qualitative differences in defect counts arise from distinct dominant dynamics at different stages: in the early steps, local defect dynamics prevail, while in the later steps the variation in stripe period becomes the primary factor, as previously reported [shimizu2023].
@@
Figure 10: Number of defects averaged over six trials. Green illustrate the count of junctions, and blue show the count of terminals.
@@ IV-C Medial axis length between defects
The second factor is the correlation between defects. In Ref. [shimizu2023], it was reported that strong correlations develop between junctions and terminals, and around step 12, where the system undergoes a transition from the quenched to the annealed state, the annihilation of junction-terminal pairs contributes to an increase in the coherence of the system. Since short junction–terminal pairs are preferentially annihilated, the mean length of the remaining pairs can increase. This trend is especially pronounced in Type B measurements for junction–terminal pairs.
Figure 11: Average medial axis length measurements between topological defects, calculated across six trials. Green segments: Length between adjacent junctions. Yellow segments: Length between adjacent junctions and terminals.
@@ IV-D Curvature
Defects discussed above serve as an indicator of geometrical inhomogeneity in the system and of deviations from the thermal equilibrium state. For stripe patterns in particular, their dynamics are crucial for systematically understanding the demagnetization dynamics from a topological perspective. On the other hand, for the relationship between the configuration of magnetic moments and the energy, the curvature provides a more direct and convenient quantity. As noted earlier, stripe formation arises from the competition between short-range exchange interactions and long-range dipole–dipole coupling. While exchange interactions favor uniform spin alignment, they also energetically prefer straight domain walls. Consequently, regions where black–white domain boundaries bend sharply carry a higher energy cost: the larger the local curvature, the greater the energetic penalty.
Fig. 12(a) displays the border curvature of the labyrinthine pattern. Both Type A and Type B images show similar curvature patterns, though Type A has slightly higher curvature in the quenched state. For both measurements, the curvature peaks around steps 9-11, coinciding with the transition from the quenched to the annealed state. The peak in curvature around step 10 may be related to the defect pair annihilation discussed in Section IV-B. Pair annihilation of defects requires substantial local deformation of the stripe pattern, which can produce sharp spatial modulations in the vicinity; consequently, the mean curvature may increase. After the annihilation process, however, the system locally relaxes to a smoother pattern with fewer defects, leading to a subsequent decrease in the curvature.
@@ IV-E Total Length
Fig. 13 (a)-(b) illustrates the total inner and border lengths of the magnetic stripe pattern images. The graphs show a lower average length at the beginning of the annealing protocol (in the quenched state), followed by stabilization at a higher average length in the later steps (in the annealed state). This increase in total length throughout the annealing protocol can be attributed to a more efficient packaging of stripes in the annealed state due to its higher spatial order as well as the reduction of the stripe period [shimizu2023].
@@ V Conclusion
Appendix A Noise Details
TABLE V: Parameters used for synthetic degradations.
Noise Type Parameters Value
Simplex Threshold [60, 150]
Scale [200, 500]
Octaves [4, 6]
Persistence [0.4, 0.6]
Lacunarity [2, 3]
AWGN $\sigma$ [0, 120]
Photometric $\alpha$ [0.8, 1]
$\gamma$ [0.4, 1]
TABLE VI: Visual comparison of segmentations using Otsu’s method and the U-Net architecture with various data augmentation schemes (no synthetic degradations, Gaussian noise and photometric transformation, only Simplex noise, and all combined synthetic degradations).
Condition Original Otsu U-Net (no degradations) U-Net (AWGN + photometric) U-Net (Simplex) U-Net (All augmentations)
Clear [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Noisy [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Occluded [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Noisy and Occluded [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
In order to account for varying degrees of degradation, synthetic noise was generated with parameters uniformly sampled from predefined ranges (Table V). These ranges were selected through visual inspection and kept intentionally broad to ensure that visually plausible degradations are well covered across the sampled parameter space.
For the Simplex noise occlusions, the parameters scale, octaves, persistence, and lacunarity control the texture of the generated patterns, while the threshold parameter determines the size of the occluded regions. Specific visual patterns are not directly associated with individual parameter values, instead resulting from the nonlinear interaction between them. We individually varied each parameter and checked whether visually plausible occlusion degradations would still occur under the parameters joint behavior.
Simplex noise was added first, sampling values from the scale, octaves, persistence and lacunarity parameter ranges. It was scaled to the $[0,255]$ range and then applied to the patch according to:
$patch_{\mathrm{occluded}}=\min\!\left(\frac{simplex\_noise\cdot 255}{threshold},\;patch\right)$ (4)
This process was repeated twice to generate overlapping Simplex noises with different parameter values, improving robustness.
Additive white Gaussian noise was then applied once, followed by the photometric transformation. Lastly, the image patches and the segmented patches are rescaled to the $[0,1]$ range prior to training and inference.
Table VI presents a visual comparison on the effect of different augmentation noise configurations and the resulting segmentation. To verify robustness, the table compares segmentation results across four scenarios: clear regions, noisy regions, occluded regions, and regions with both noise and occlusion.
Without any data augmentation, U-Net’s segmentations visually resembled Otsu thresholding, containing both irregular boundaries and occlusions. Training the model with both AWGN and photometric transformation resulted in smoother and better-defined borders between light and dark regions, although issues with occlusion degradations persisted. When trained exclusively with Simplex noise, the segmentation was similar to Otsu’s method in clear regions, with moderate improvements in boundary smoothness and occlusion filtering. When all augmentations were used during training, segmentation borders remained smooth and well-defined, and filtering of occluded regions further improved. These results suggest that Simplex noise alone does not adequately model true occlusions found in real magnetic labyrinthine images, leading to limited generalization to real occlusions.
Appendix B Validation Metrics
Validation performance metrics are presented in table. VII. An ablation study on the impact of different synthetic degradations in training is provided. Using all AWGN, photometric transformation and Simplex noise yields the best overall balance between overlap, boundary distance, and length deviations.
Fig. 14 showcase the validation loss curve of the U-Net model using all augmentations, while Fig. 15 showcase the validation performance metrics throughout the optimization process.
TABLE VII: Validation performance metrics.
Model F1 score ($\uparrow$) IoU ($\uparrow$) ASSD ($\downarrow$) $\DeltaInner Δ\DeltaBorder$
Otsu 0.944 0.895 1.14 3.33% 3.43%
SegFormer-B3 0.951 0.907 1.06 -0.76% -0.93%
U-Net (all augmentations) 0.953 0.910 1.01 -0.57% -0.10%
No augmentation 0.940 0.888 1.24 3.33% 4.79%
AWGN + photometric 0.947 0.900 1.12 0.01% 0.59%
Simplex only 0.943 0.893 1.19 1.19% 2.29%
Figure 14: Validation loss curve.
Figure 15: Validation curves.
[…]